OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 1/x , g(x) = |x|-2 What is the domain of (f º g) (x)?

OpenStudy (anonymous):

1st proceed by ur own...what will u do?

OpenStudy (anonymous):

honestly i have no clue.

OpenStudy (anonymous):

and i only have a few minutes to do this.

OpenStudy (anonymous):

ok 1st u have to find the composite function..f(g(x))

OpenStudy (anonymous):

so g(x)= -x+2, x<0 = x+2 , x> or = 0 f(x)= 1/x, x not equal to zero.

OpenStudy (anonymous):

is it {x|x ≠ 0}

OpenStudy (anonymous):

yes for f(x)=1/x

OpenStudy (anonymous):

or {x|x ≠ -2, x ≠ 2}

OpenStudy (anonymous):

u have written g(x)= I x I -2

OpenStudy (anonymous):

{x|x ≠ -2, x ≠ 2} ...from where did u get it?

OpenStudy (anonymous):

Theresa are the options: A: {x | x e reall numbers } B: {x | x ≠ 0} C: {x |x ≠ -2, x ≠2} D: {x | x ≠ 0, x ≠ 2}

OpenStudy (anonymous):

oh...u were asking answers directly...i was tell u ...how to solve this?

OpenStudy (anonymous):

yes answer is B

OpenStudy (anonymous):

ok